OpenStudy (anonymous):

What is the value of f(5) for the function f(x) = 3x - 4?

OpenStudy (anonymous):

f(5)= 3(5)-4 f(5)= 15-4 f(5)= 11

OpenStudy (anonymous):

Thank you! I suck at functions..

OpenStudy (anonymous):

Would you mind doing another one with me?

OpenStudy (debbieg):

f(some number) just means to plug in that number wherever you see "x", and simplify. That will give you the function value, "evaluated at x". @kdekle showed you how to do this above. Why not try your next problem yourself, and then just post what you did, and someone can tell you if you're right? I bet you can do it! :)

OpenStudy (anonymous):

It has two letters now though, and i dont know where to begin. What is the value of f[g(5)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

Does x just equal five in both of them?

OpenStudy (anonymous):

no you start from the inside and work out. so the inside says g(5). so Work out 2(5)-5. Now THAT answer is the x value in f.

OpenStudy (anonymous):

OH! (:

OpenStudy (debbieg):

That one looks a little bit different... but it isn't really. :) Just do it in two steps. 1. find g(5) (just like how you did f(5) above, except you'll use g(x) and plug in 5 for x) 2. Once you have g(5), which is just going to be some number, plug THAT number into f(x) for x and proceed as before! :)

OpenStudy (anonymous):

2(5)-5 10-5 5 2(5)+1 10+1 11 ?

OpenStudy (anonymous):

exactly!

OpenStudy (debbieg):

Remember, when you see f(something) (or g(something) or t(something) or Q(something).... the point being, it does not matter what letter is used to "name" the function, that's just a name)..... what it means it to take the "rule" that is given to define f, and plug in the (something) for the variable. So in this case, where you have f[g(5)], the "something" that is to go into f, is g(5). So first you need to find g(5), which you do by using the rule for g(x), with x=5. :)

OpenStudy (debbieg):

You're welcome, and no problem, @kdekle got you started and I just gave an "assist", lol. :)